Solve:
8(2f-3)= 4(4f-8)

Answer by rmromero(383)


Solve:
8(2f-3)= 4(4f-8)
Use distributive property to simplify both sides.
8(2f) - 8(3) = 4(4f) - 4(8)
16f - 24 = 16f - 32

16f - 16f - 24 = 16f - 16f - 32
-24 = -32 Not True
Since -24 is not equal to -32, therefore there is no solution.